find the lengths of the apothem and the radius of a square whose sides have length 10in ?
Answer by jim_thompson5910(35256)

For any square, the apothem is always half as long as the side length

apothem = (1/2)*(side length)
apothem = (1/2)*(10)
apothem = 5

The apothem is 5 units long. The apothem is the segment starting at the center of the regular polygon and it stretches to the midpoint of any side of the regular polygon.

-------------------------------------

If you draw out the square, add in the apothem and radius, you'll see a right triangle form.

This right triangle has legs of 5 and 5. The hypotenuse is unknown. Let's call it x for now

Pythagorean theorem:
a^2 + b^2 = c^2
5^2 + 5^2 = x^2
25 + 25 = x^2
50 = x^2
x^2 = 50
x = sqrt(50)
x = sqrt(25*2)
x = sqrt(25)*sqrt(2)
x = 5*sqrt(2)

So the radius of the square is 5*sqrt(2)
The radius of a regular polygon (aka the circumradius) is the segment going from the center to any vertex of the regular polygon.

-----------------------------------

Answers:

Apothem = 5 inches
(circum)radius = 5*sqrt(2) inches